OpenStudy (anonymous):

Use implicit differentiation to find dy/dx where x^2y = 1 + e^y.

OpenStudy (anonymous):

i can get as far as... 2xy * y(dy/dx) then i don't know how to do the e^y

hartnn (hartnn):

use chain rule, d/dx(e^y) = e^y d/dx(y)

hartnn (hartnn):

which is e^y y' got this ?

OpenStudy (anonymous):

hmmm nope. i don't see it...

hartnn (hartnn):

d/dx of e^x is just e^x but here y is the function of x so u need chain rule

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = \frac{ -2xy }{ x^2}\] that's as far as i can get...

OpenStudy (anonymous):

that whole e to the y throws me off so much...

hartnn (hartnn):

no, u need to take care of e^y also and \(x^2y\) needs a product rule

OpenStudy (anonymous):

kira that's the wrong questoin...

OpenStudy (anonymous):

i know i do. i'm just saying i don't know how to take care of e^y

hartnn (hartnn):

and that i said e^y needs chain rule because its function of x d/dx (e^y) = e^y d/dx(y) = e^y dy/dx did u get this ? if not which step ?

OpenStudy (anonymous):

no that just confuses me... can you give another example?

OpenStudy (kira_yamato):

In that case, d/dx (e^y) = d/dy (e^y) * dy/dx = e^y * dy/dx

OpenStudy (anonymous):

yeah that's what hartnn said...

hartnn (hartnn):

sure d/dx (sin y) = cos y d/dx(y) = cos y dy/dx

OpenStudy (kira_yamato):

If z = ln y and we want to find dz/dx dz/dx = 1/y * dy/dx

OpenStudy (kira_yamato):

dz/dx = dz/dy * dy/dx

hartnn (hartnn):

u got two examples :)

OpenStudy (anonymous):

well i haven't used cos or sin yet so that's a little confusing... let me see...

hartnn (hartnn):

d/dx( y^3) = 3y^2 d/dx (y) = 3y^2 dy/dx

hartnn (hartnn):

want more examples ?

OpenStudy (anonymous):

sure i'll take one more. i do get it with number exponents. like 2y^2 would be 4y(dy/dx) right?

hartnn (hartnn):

yes, absolutely correct :)

OpenStudy (anonymous):

then what is up with e^y?! :(

OpenStudy (anonymous):

i like your avatar by the way

hartnn (hartnn):

first , whats d/dx of e^x

hartnn (hartnn):

thanks :)

OpenStudy (anonymous):

e^x

hartnn (hartnn):

right , but now here its not e^x , its e^y where y is the function of x so using chain rule, we have to differentiate y after we differentiate e^y so we get like this d/dx (e^y) = e^y . d/dx (y)

OpenStudy (anonymous):

...okay...

hartnn (hartnn):

so after diff. u get 2xy + x^2 dy/dx = e^y dy/dx got this ?

hartnn (hartnn):

now isolate dy/dx

OpenStudy (anonymous):

i know i'll have 2xy over x^2 on the right siide. and dy/dx on the left, but i don't know what to do with e^y(dy/dx) that's on the right...

OpenStudy (anonymous):

*negative 2xy

hartnn (hartnn):

2xy + x^2 dy/dx = e^y dy/dx x^2 dy/dx - e^y dy/dx = -2xy (x^2 - e^y ) dy/dx = -2xy can u finish it off ?

OpenStudy (anonymous):

\[\frac{ -2xy }{ x^2 - e^y }\]

OpenStudy (anonymous):

what confused me was the double dy/dx... i didn't think of it like factoring.

hartnn (hartnn):

thats correct :)

OpenStudy (anonymous):

damn you math!! but thank you hartnn!!

hartnn (hartnn):

welcome ^_^